OpenStudy (anonymous):

I really need help with this one Find the slope of the line containing the given points (-8/7, -9/10) and (7/5, -9/10)

OpenStudy (anonymous):

slope = (y2-y1)/(x2-x1) (x1, y1)= (-8/7, -9/10) (x2,y2)=(7/5, -9/10)

OpenStudy (anonymous):

i know the formula i have just never done it with fractions before, thats the part i need help with

OpenStudy (anonymous):

Wait, the y values in both points are the same. That means the line is horizontal.

OpenStudy (anonymous):

That means the slope is 0

OpenStudy (anonymous):

oh i was think it was undefined but i remember now that's only when it vertical thanks

OpenStudy (anonymous):

Yep!